# Fairness-in-the-Workflow: How Machine Learning Practitioners at Big Tech Companies Approach Fairness in Recommender Systems

### v2

@@ Abstract.
Recommender systems (RS), which are widely deployed across high-stakes domains, are susceptible to biases that can cause large-scale societal impacts. Researchers have proposed methods to measure and mitigate such biases—but translating academic theory into practice is inherently challenging. RS practitioners must balance the competing interests of diverse stakeholders, including providers and users, and operate in dynamic environments. Through a semi-structured interview study (N=11), we map the RS practitioner workflow within large technology companies, focusing on how technical teams consider fairness internally and in collaboration with other (legal, data, and fairness) teams. We identify key challenges to incorporating fairness into existing RS workflows: defining fairness in RS contexts, particularly when navigating multi-stakeholder and dynamic fairness considerations. We also identify key organization-wide challenges: making time for fairness work and facilitating cross-team communication. Finally, we offer actionable recommendations for the RS community, including HCI researchers and practitioners.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainccs: Information systems Recommender systemsccs: Social and professional topics User characteristicsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recommender systems (RS) are inherently social and collaborative entities (Stray et al., 2024) which connect providers (e.g., sellers, content creators) with users (e.g., buyers, content consumers). RS are widely deployed across high-stakes domains including e-commerce (Gretzel and Fesenmaier, 2006; Zalmout et al., 2021) and social media (Lada et al., 2021; Belli et al., 2020; Anandhan et al., 2018)—but they are susceptible to biases that can cause large-scale psychological, ideological, and societal impacts within those domains (Li et al., 2021b; Chen et al., 2023; Kay et al., 2015; Sweeney, 2013). For example, Amazon’s e-commerce RS shows preference to Amazon’s own products even when competing products have superior reviews, resulting in an effective monopoly (Tate, 2023), and X’s social media RS disproportionately amplifies right-leaning political content, resulting in potential electoral impacts (Ye et al., 2025; Huszár et al., 2022). ML, HCI, and policy researchers have proposed a variety of methods to measure and mitigate such biases in order to improve fairness in RS (Feng and Shah, 2022; Koenecke et al., 2023; Smith et al., 2024; Abdollahpouri et al., 2019). However, appropriately translating any academic theory into industry application is inherently challenging because it requires navigating real-world constraints and balancing conflicting stakeholder interests (Holstein et al., 2019; Deng et al., 2022; Madaio et al., 2020; Beutel et al., 2019b; Veale et al., 2018; Metcalf et al., 2019; Robertson et al., 2022; Lee and Singh, 2021; Quiñonero Candela et al., 2023; Rieke et al., 2022; Smith et al., 2022). Incorporating fairness considerations into RS development is particularly challenging because RS operate within a landscape where fairness is not a static concept but a dynamic one, with meanings that shift across different contexts (Smith et al., 2024, 2025; Ekstrand et al., 2022; Smith et al., 2023b).
@@
The complexity of RS is amplified by its operation in a multi-sided market, involving diverse stakeholders—both providers and end-users, each with heterogeneous make-ups (Burke, 2017; Abdollahpouri and Burke, 2019). RS practitioners who seek to incorporate fairness into their workflows therefore face the unique challenge of balancing multiple competing sets of stakeholder interests. Furthermore, RS practitioners operate subject to multiple ethical and regulatory constraints, such as the need to ensure the dissemination of trustworthy information, and sometimes must make tradeoffs between prioritizing fairness and those other considerations. RS practitioners also do not operate in a static environment. Their roles require not only identifying existing sources of unfairness, but anticipating emerging ones and mitigating their long-term effects, which can be exacerbated by feedback loops (Burke, 2017; Li et al., 2021b). This task demands a high level of foresight and adaptability—practitioners must be not only reactive, but also proactive in their approach to improving fairness in RS.
@@
The ubiquity of RS today underscores the critical role that RS practitioners play in shaping society’s online interactions. Technical RS teams serve as the central hub where innovation, ethical considerations, and practical applications intersect. Here, the core challenges of fairness in RS—such as conflicting stakeholder interests and uncertain long-term impacts—are actively addressed and navigated. Hence, building on existing research, we explore the unique nuances of how machine learning (ML) practitioners at ‘big tech’ companies approach fairness in RS. In particular, we map the RS practitioner workflow, with a particular emphasis on how technical teams attempt to incorporate fairness (a) into their existing internal practices and (b) in collaboration with legal, data, and fairness teams. This allows us to understand how technical RS practitioners approach fairness in RS—a focus that is critical in understanding how to bridge academic theory on fairness in RS with actual practices. To do this, we conduct semi-structured interviews with RS technical practitioners who are integral members of ML teams at big tech companies (N=11). We examine what they have done to integrate fairness into their workflows, as well as the technical and organizational challenges they have faced to doing so. Technically, practitioners struggle to define fairness in the context of RS, especially because they must account for complexities inherent to RS, which include considering fairness across large numbers of (conflicting) stakeholder groups and working in dynamic settings with deployed systems that are susceptible to feedback loops. Organizationally, practitioners struggle with a lack of time to conduct fairness work and with developing a ‘fairness lingua franca’ that allows for clear communication across multidisciplinary teams. Our findings reveal opportunities for practitioners, ML researchers, and HCI researchers to improve fairness in RS by improving collaboration practices across industry teams and between industry and academia. Our key contributions are:
@@
  • Actionable steps towards addressing the challenges we identify by improving collaboration, communication, and documentation practices so that RS practitioners are better equipped to take lessons not only from industry fairness teams but also from academic researchers
We proceed as follows. First, we provide an overview of prior work on fairness considerations for RS, as well as prior work on the challenges inherent to translating academic theory into practice (§2). We describe our semi-structured interview and data analysis approach (§3). Then, we provide an overview of the RS workflow (§4). We describe technical challenges (§5) and organizational challenges (§6) to implementing fairness in RS workflows. We close with calls to action for the RS, ML fairness, and HCI communities (§7) as well as a discussion of the limitations of our work and potential paths forward (§8).
@@ 2. Related Work
Here, we provide a brief overview of the academic literature on ML fairness, with a particular focus on fairness for RS. Then, we discuss prior work on implementing fairness in deployed systems, with a focus on the challenges inherent to translating academic principles into practice and on the implications of ML fairness for industries and organizations.
2.1. Bias and Fairness in Machine Learning
Fairness and bias are closely related but distinct. Bias refers to systematic errors or deviations in data or models producing skewed outcomes for certain groups (Hutchinson and Mitchell, 2019). Fairness, by contrast, concerns the ethical acceptability of those biased outcomes (Corbett-Davies et al., 2023). We focus on fairness in this paper. [^1] We note that during interviews, interview participants sometimes used the terms ‘bias’ to refer to what we generally call ‘unfairness,’ or ‘debiasing’ to refer to what we generally call ‘improving fairness,’ so we use those terms when quoting participants directly. Because formal fairness definitions can be mutually incompatible, organizations must choose which principles to prioritize (Hutchinson and Mitchell, 2019; Ekstrand et al., 2022). For example, in a video-sharing platform, while one approach to fairness might involve using straightforward viewer metrics for ranking videos, an alternative approach might incorporate randomness in search rankings to counteract the “rich-get-richer” effect (Chen et al., 2023) wherein top-ranked videos get disproportionately more future views. The choice of how to define fairness often intersects with legal obligations (e.g., anti-discrimination laws) and business goals (Hutchinson and Mitchell, 2019; Ali et al., 2023). Our study builds on the fair ML literature by examining how fairness is approached in practice in RS workflows.
2.2. Fairness for Recommender Systems
RS pose unique fairness challenges due to multi-sided stakeholder needs (Chen et al., 2023; Wang et al., 2024b; Abdollahpouri and Burke, 2019; Burke, 2017). RS practitioners must balance the (potentially competing) interests of providers (e.g., sellers) and end-users (e.g., buyers)  (Burke et al., 2018; Singh and Joachims, 2018; Sonboli et al., 2021; Leonhardt et al., 2018; Li et al., 2021a; Stray et al., 2024; Mehrotra et al., 2018). For instance, a user-centric fairness approach might target equally accurate recommendations across demographics (Sonboli et al., 2021), while a provider-focused approach might seek equitable exposure for underrepresented creators (Abdollahpouri et al., 2019)—and these two interests can easily be in direct conflict with one another (Hutchinson and Mitchell, 2019; Sonboli et al., 2021; Ekstrand et al., 2022; Smith et al., 2023b). A system that prioritizes promoting content from historically marginalized providers might inadvertently serve users content that is misaligned with their previously stated interests (Smith et al., 2023b). And as Smith et al. (2024) point out, such a system can even have potentially negative effects for providers by exposing them to harassment from users who are outside of their preferred target audience. Complicating things further, a single platform may contain multiple types of providers, such as social media content creators and paid advertisers, whose interests may conflict as well. Despite the fact that RS are a commonly used technology, much ML fairness research has focused on classification problems (Barocas and Selbst, 2016; Chouldechova, 2017; Kleinberg et al., 2017; Dwork et al., 2012; Feldman et al., 2015; Zafar et al., 2017a; Corbett-Davies et al., 2017, 2023; Friedler et al., 2019; Kilbertus et al., 2017; Hardt et al., 2016b; Pleiss et al., 2017; Kamiran and Calders, 2009; Buolamwini and Gebru, 2018; Obermeyer et al., 2019; Friedler et al., 2021; Lee et al., 2024; Grgic-Hlaca et al., 2016; Hardt et al., 2016a; Sharifi-Malvajerdi et al., 2019; Zafar et al., 2017b; Borgs et al., 2019; Tal, 2023)—meaning that many existing fairness objectives and metrics cannot be directly applied to RS. Large-scale RS also raise unique challenges in practice related to sparse user-content interaction data, because users may only interact with a tiny subset of overall content (Li et al., 2021a; Kearns et al., 2019; Beutel et al., 2019b, a, 2018; Lum et al., 2022). Finally, as RS use user-content interaction data generated from the recommendation process to inform their future results, they risk feedback loops that amplify disparities (Keller, 2021). Our study contributes perspectives on implementing ML fairness in practice that are specific to RS, taking these unique challenges into account.
@@ 2.3. Fairness in Practice
In practice, fairness efforts do not occur in isolation. Rather, they interact—and may come into tension—with organizational priorities, such as maximizing profits. Acquiring organizational support to conduct responsible AI initiatives often requires linking fairness to explicit business or product goals (Scheuerman, 2024; Ali et al., 2023; Garcia-Gathright et al., 2018; Rakova et al., 2021; Veale et al., 2018; Wang et al., 2024a). Legal teams can also play a role in bringing about responsible AI efforts, although they often focus on privacy or liability rather than proactive fairness (Veale and Binns, 2017). However, regulations often vary across legal jurisdictions, complicating fairness efforts for multinational platforms that use RS. For example, RS in the EU are subject to the Digital Services Act (DSA), which authorizes authorities to enforce yearly assessments of online platforms regarding fairness issues, illegal content, and disruption to human rights (Madiega, 2020). The DSA does not apply to RS in the US (Fiordalisi, 2025); however, individual states have created their own fairness regulations, such as California’s 2022 state law AB 587, [^2] https://openstates.org/ca/bills/20212022/AB587/ which aims to provide transparency about large platform recommendations to avoid disinformation and hate speech. When fairness standards are not mandated by law, legal teams may actually hinder fairness efforts due to concerns that they could add to organizational liability (Costanza-Chock et al., 2022; Wright et al., 2024). If institutional support is lacking, employees may choose to “individualize risk” (Ali et al., 2023) by taking personal responsibility for incorporating fairness into their workflows. Prior work has repeatedly shown that these efforts can be challenging, as practitioners struggle to reconcile incompatible fairness metrics (Madaio et al., 2022; Moharana et al., 2025; Smith et al., 2025), work with messy real-world data (Cramer et al., 2019; Holstein et al., 2019; Rieke et al., 2022; Guerdan et al., 2025; Passi and Barocas, 2019; Sambasivan et al., 2021), and adapt general-purpose academic tools and metrics to their specific real-world needs (Deng et al., 2022; Harvey et al., 2025, 2024; Lee and Singh, 2021; Balayn et al., 2023; Kaur et al., 2020; Richardson et al., 2021; Ojewale et al., 2025). Many of these challenges are exacerbated by organizational challenges, such as communicating and collaborating across different teams and within cross-functional teams (Deng et al., 2023b). Researchers including Madaio et al. (2020), Deng et al. (2023a), and Costanza-Chock et al. (2022) have repeatedly found that a particular pain point when attempting to improve algorithmic fairness in practice is stakeholder involvement: practitioners struggle to involve the perspectives of end users in fairness assessments and improvements.
@@
With the notable exception of Smith et al. (Smith et al., 2023b, a, 2024, 2025), most prior work on implementing fairness in practice has focused broadly on general-purpose ML or specifically on fairness in classification models, NLP models, or other non-RS applications of ML. However, we expect the issues raised by prior work to be even more pressing in the RS space, where fewer academic toolkits are available to scaffold fairness efforts, and where the interests of multiple stakeholders (providers and users) may be in direct competition. In fact, Smith et al. have previously identified tensions between competing stakeholders through interview studies with individuals who interact with RS but do not work directly on technical RS teams, including providers (content creators and dating app users) (Smith et al., 2024) and ML fairness experts (Smith et al., 2025). Most similar to our work, Smith et al. have also conducted interview studies with RS practitioners themselves: employees of an anonymous “commercial consumer-provider recommender system” (Smith et al., 2023a) and of a microlending platform (Smith et al., 2023b). Both studies explored challenges related to selecting fairness metrics in the context of RS, again identifying competing stakeholder interests as a key challenge (Smith et al., 2023b, a). In contrast, we interview technical RS practitioners across multiple big tech companies, and look beyond fairness metrics to how fairness is incorporated into RS workflows more broadly. In summary, we build on the extensive body of prior work by focusing explicitly on the RS workflow, with a particular spotlight not only on how RS practitioners across multiple big tech companies face technical challenges in the face of RS-specific complexity, but also how they navigate organizational challenges stemming from cross-team interactions with legal, data, and fairness practitioners.

### v3

# arXiv:2505.19441v3[cs.HC] 27 Feb 2026

@@ Abstract.
Recommender systems (RS), which are widely deployed across high-stakes domains, are susceptible to biases that can cause large-scale societal impacts. Researchers have proposed methods to measure and mitigate such biases—but translating academic theory into practice is inherently challenging. Through a semi-structured interview study (N=11), we map the RS practitioner workflow within large technology companies, focusing on how technical teams consider fairness internally and in collaboration with legal, data, and fairness teams. We identify key challenges to incorporating fairness into existing RS workflows: defining fairness in RS contexts, balancing multi-stakeholder interests, and navigating dynamic environments. We also identify key organization-wide challenges: making time for fairness work and facilitating cross-team communication. Finally, we offer actionable recommendations for the RS community, including practitioners and HCI researchers.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791347isbn: 979-8-4007-2278-3/2026/04ccs: Information systems Recommender systemsccs: Social and professional topics User characteristicsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recommender systems (RS) are inherently social and collaborative entities (Stray et al., 2024; Li et al., 2025) which connect providers (e.g., sellers, content creators) with users (e.g., buyers, content consumers). RS are widely deployed across high-stakes domains including e-commerce (Gretzel and Fesenmaier, 2006; Zalmout et al., 2021) and social media (Lada et al., 2021; Belli et al., 2020; Anandhan et al., 2018)—but they are susceptible to biases that can cause large-scale psychological, ideological, and societal impacts (Li et al., 2021b; Chen et al., 2023; Kay et al., 2015; Sweeney, 2013). For example, Amazon’s e-commerce RS shows preference to Amazon’s own products even when competing products have superior reviews, resulting in an effective monopoly (Tate, 2023), and X’s social media RS disproportionately amplifies right-leaning political content, resulting in potential electoral impacts (Ye et al., 2025; Huszár et al., 2022). ML, HCI, and policy researchers have proposed a variety of methods to measure and mitigate such biases in order to improve fairness in RS (Feng and Shah, 2022; Koenecke et al., 2023; Smith et al., 2024; Abdollahpouri et al., 2019). However, appropriately translating any academic theory into industry application is inherently challenging because it requires navigating real-world constraints and balancing conflicting stakeholder interests (Holstein et al., 2019; Deng et al., 2022; Madaio et al., 2020; Beutel et al., 2019b; Veale et al., 2018; Metcalf et al., 2019; Robertson et al., 2022; Lee and Singh, 2021; Quiñonero Candela et al., 2023; Rieke et al., 2022; Smith et al., 2022). Incorporating fairness considerations into RS development is particularly challenging because RS operate within a landscape where fairness is not a static concept but a dynamic one, with meanings that shift across different contexts (Smith et al., 2024, 2025; Ekstrand et al., 2022; Smith et al., 2023b).
@@
The complexity of RS is amplified by its operation in a multi-sided market, involving diverse stakeholders—both providers and end-users, each with heterogeneous make-ups (Burke, 2017; Abdollahpouri and Burke, 2019). RS practitioners who seek to incorporate fairness into their workflows therefore face the unique challenge of balancing multiple competing sets of stakeholder interests. Furthermore, RS practitioners operate subject to multiple ethical and regulatory constraints, such as the need to ensure the dissemination of trustworthy information, and sometimes must make tradeoffs between prioritizing fairness and those other considerations. RS practitioners also do not operate in a static environment. Their roles require not only identifying existing sources of unfairness, but anticipating emerging ones and mitigating their long-term effects, which can be exacerbated by user feedback loops (Burke, 2017; Li et al., 2021b, 2025). This task demands a high level of foresight and adaptability—practitioners must be proactive in their approach to improving fairness in RS.
@@
The ubiquity of RS today underscores the critical role that RS practitioners play in shaping society’s online interactions. Technical RS teams serve as the central hub where innovation, ethical considerations, and practical applications intersect. Here, the core challenges of fairness in RS are actively addressed and navigated. Hence, building on existing research, we explore the unique nuances of how machine learning (ML) practitioners at ‘big tech’ companies approach fairness in RS. In particular, we map the RS practitioner workflow, with an emphasis on how technical teams attempt to incorporate fairness (a) into their existing internal practices and (b) in collaboration with legal, data, and fairness teams. This allows us to understand how technical practitioners approach fairness in RS—a focus that is critical in understanding how to bridge academic theory with actual practices. To do this, we conduct semi-structured interviews with technical RS practitioners who are integral members of ML teams at big tech companies (N=11). We examine what they have done to integrate fairness into their workflows, as well as the technical and organizational challenges they have faced to doing so. Technically, practitioners struggle to define fairness in the context of RS, especially because they must account for complexities inherent to RS, which include considering fairness across large numbers of (conflicting) stakeholder groups and working in dynamic settings with deployed systems that are susceptible to feedback loops. Organizationally, practitioners struggle with a lack of time to conduct fairness work and with developing a ‘fairness lingua franca’ that allows for clear communication across multidisciplinary teams. Our findings reveal opportunities for practitioners and HCI researchers to improve fairness in RS. Our key contributions are:
@@
  • A set of actionable recommendations to better integrate fairness into RS workflows, highlighting both where big tech companies and practitioners are responsible for implementing evidence-based changes as well as where novel HCI research can support real-world fairness efforts
We proceed as follows. First, we provide an overview of prior work on fairness considerations for RS and on the challenges inherent to translating academic theory into practice (§2). We describe our semi-structured interview and analysis approach (§3). Then, we provide an overview of the RS workflow (§4). We describe technical challenges (§5) and organizational challenges (§6) to implementing fairness in RS workflows. We make calls to action for RS practitioners and HCI researchers (§7) and close with a discussion of the limitations of our work and potential paths forward (§8).
@@ 2. Related Work
Here, we provide a brief overview of the academic literature on ML fairness, with a focus on fairness for RS. We also discuss prior work on implementing fairness in deployed systems, with a focus on the challenges inherent to translating academic principles into practice.
2.1. Fairness in Machine Learning
Fairness and bias are closely related but distinct. Bias refers to systematic errors or deviations in data or models that produce skewed outcomes for certain groups (Hutchinson and Mitchell, 2019). Fairness, by contrast, concerns the ethical acceptability of those biased outcomes (Corbett-Davies et al., 2023). We focus on fairness in this paper. [^1] We note that during interviews, participants sometimes used the terms ‘bias’ to refer to what we generally call ‘unfairness,’ or ‘debiasing’ to refer to what we generally call ‘improving fairness,’ so we use those terms when quoting participants directly. Fairness is an essentially contested concept, meaning that there is no single agreed-upon approach for defining fairness (Jacobs and Wallach, 2021). A dominant fairness paradigm focuses on the outcomes of technical systems: systems are group fair to the extent that they produce similar outcomes for different groups of individuals, and individually fair to the extent that they treat ‘similar’ individuals similarly (Dwork et al., 2012; Barocas and Selbst, 2016). Although these approaches sound simple, they are anything but: group fairness can be defined using multiple metrics, many of which are impossible to simultaneously satisfy, and individual fairness requires a way of determining how similar any two individuals are, which is often infeasible in practice (Chouldechova, 2017; Kleinberg et al., 2017; Zemel et al., 2013). An alternative paradigm draws on the theory of procedural justice (Rawls, 1999), arguing that systems are fair to the extent that they follow fair processes—although what makes a process fair is not settled, and may involve anything from only using certain features (Grgic-Hlaca et al., 2016) to incorporating multiple stakeholder groups in participatory system development (Lee et al., 2019). Finally, fairness is but one component of the broader field of ethical or responsible AI. Thus, the choice of how to define fairness often intersects with other considerations. These may include legal obligations (e.g., anti-discrimination laws) (Hildebrandt, 2022; Magrani and da Silva, 2024), explainability (Kunkel et al., 2019), user privacy considerations (Sun et al., 2023), and the extent to which systems facilitate appeal and recourse of algorithmic decisions (Jin and Salehi, 2024). Ultimately, organizations must choose which principles to prioritize, and the choice is far from simple (Hutchinson and Mitchell, 2019; Ekstrand et al., 2022).
Fairness for Recommender Systems
Defining fairness is particularly challenging in the context of RS because much ML fairness research has focused on classification problems, meaning that many existing fairness objectives and metrics may not be directly applicable to RS. For example, although fairness in classification problems often focuses on outcomes, what an outcome means in the context of RS is not necessarily clear: outcomes can be thought of in terms of not only allocations (e.g., whose content gets recommended) but also representations (e.g., whether recommended content contributes to the stereotyping, erasure, or demeaning of social groups) (Blodgett et al., 2020). Nevertheless, researchers have proposed a variety of approaches for measuring and improving fairness in recommendations and ranking (Zehlike et al., 2017; Asudeh et al., 2019; Biega et al., 2018; Ali et al., 2019; Beutel et al., 2019a; Diaz et al., 2020; Ekstrand et al., 2018; Geyik et al., 2019; Li et al., 2021a; Zehlike and Castillo, 2020; Singh and Joachims, 2018; Steck, 2018; Yang and Stoyanovich, 2017; Rastogi and Joachims, 2024). [^2] For comprehensive overviews of fairness in RS, we direct the reader to Ekstrand et al. (2022), Patro et al. (2022), and Zehlike et al. (2022a, b). For a survey of fair ranking metrics, we point to Smith et al. (2023a). As discussed above, defining fairness (e.g., by choosing a fairness metric) is challenging in ML generally. It is uniquely challenging in the context of RS because RS require balancing the (potentially competing) interests of multiple stakeholders who are dynamically interacting with each other and the RS platform at massive scales (Chen et al., 2023; Wang et al., 2024b; Abdollahpouri and Burke, 2019; Burke, 2017). In particular, RS practitioners must consider the interests of providers (e.g., sellers) and end-users (e.g., buyers)  (Burke et al., 2018; Singh and Joachims, 2018; Sonboli et al., 2021; Leonhardt et al., 2018; Li et al., 2021a; Stray et al., 2024; Mehrotra et al., 2018). For instance, a user-centric fairness approach might target equally accurate recommendations across demographics (Sonboli et al., 2021), while a provider-focused approach might seek equitable exposure for underrepresented creators (Abdollahpouri et al., 2019)—and these two interests can easily be in direct conflict with one another (Hutchinson and Mitchell, 2019; Sonboli et al., 2021; Ekstrand et al., 2022; Smith et al., 2023b). A system that prioritizes promoting content from historically marginalized providers might inadvertently serve users content that is misaligned with their previously stated interests (Smith et al., 2023b). In fact, recent user studies have found that users often prefer gender-biased recommendations over ‘debiased’ ones, even when they are made aware that bias is present (Wang et al., 2023; Krause et al., 2025). As Smith et al. (2024) point out, a ‘debiased’ system can even have potentially negative effects for providers by exposing them to harassment from users who are outside of their preferred target audience. Complicating things further, a single platform may contain multiple types of providers, such as social media content creators and paid advertisers, whose interests may conflict as well. Even within a stakeholder group, how fairness should be defined is not straightforward. For example, when considering providers in a video-sharing platform, while one approach to fairness might involve rewarding providers for producing high-quality content by using straightforward viewer metrics for ranking videos, an alternative approach might incorporate randomness in search rankings to counteract the “rich-get-richer” effect (Chen et al., 2023) wherein top-ranked videos get disproportionately more future views.
The massive scale of RS also raises unique challenges in practice. As Anthis et al. (2025) argue, technologies that are used by large populations across many contexts, like RS, simply cannot be made fair for every context in which they operate and every population subgroup with which they interact. Scale also causes technical challenges related to sparse user-content interaction data, because users may only interact with a tiny subset of overall content (Li et al., 2021a; Kearns et al., 2019; Beutel et al., 2019b, a, 2018; Lum et al., 2022). Finally, as RS use user-content interaction data generated from the recommendation process to inform their future results, they risk feedback loops that amplify disparities (Keller, 2021; Li et al., 2025). Feedback loops themselves are multi-faceted: users can provide implicit feedback (e.g., quickly scrolling past a video) or explicit feedback (e.g., requesting not to see a video again) (Oard et al., 1998; Li et al., 2025). Users provide explicit feedback less consistently than they provide implicit feedback (Lustig et al., 2022); however, implicit feedback is harder to interpret and is often itself shaped by users’ folk theories of how the RS they interact with work (Eslami et al., 2016; DeVito et al., 2017; Xiao et al., 2025). The choices by RS practitioners about what feedback to include—and the choices by users about how to convey and organize their feedback—can thus have major impacts on fairness in RS. Our study builds on the fair ML literature by examining how inherent fairness tensions are approached in practice and the extent to which academic definitions of fairness guide industry approaches, taking the unique challenges of RS fairness into account.
@@ 2.2. Fairness in Practice
In practice, fairness efforts do not occur in isolation. Rather, they interact—and may come into tension—with organizational priorities, such as maximizing profits. Acquiring organizational support to conduct responsible AI initiatives often requires linking fairness to explicit business or product goals (Scheuerman, 2024; Ali et al., 2023; Garcia-Gathright et al., 2018; Rakova et al., 2021; Veale et al., 2018; Wang et al., 2024a). Regulation can also play a role in bringing about responsible AI efforts. However, regulations often vary across legal jurisdictions, complicating fairness efforts for multinational platforms that use RS. For example, RS in the EU are subject to the Digital Services Act (DSA), which authorizes yearly assessments of online platforms regarding fairness issues, illegal content, and disruption to human rights (Madiega, 2020). The DSA does not apply to RS in the US (Fiordalisi, 2025); however, individual states have created their own fairness regulations, such as California’s 2022 state law AB 587, [^3] https://openstates.org/ca/bills/20212022/AB587/ which aims to provide transparency about large social media RS to avoid disinformation and hate speech. When fairness standards are not mandated by law, organizations may avoid assessing fairness altogether due to concerns that uncovering unfairness could add to organizational liability (Costanza-Chock et al., 2022; Wright et al., 2024). If institutional support is lacking, employees may choose to “individualize risk” (Ali et al., 2023) by taking personal responsibility for incorporating fairness into their workflows (Ganesh et al., 2025).
Operationalizing Fairness
However efforts are initiated, doing fairness in practice requires adhering to a specific operationalization of fairness. For example, documentation efforts like datasheets for datasets (Gebru et al., 2021), model cards for model reporting (Mitchell et al., 2019), and FactSheets (Arnold et al., 2019) call on practitioners to publicly share standardized reports describing the intended use, disaggregated performance, and other fairness-related characteristics of their data and models. Similarly, fairness checklists, which are inspired by safety checklists used in medicine and aviation, are intended to encourage practitioners to consider fairness issues and mitigate potential sources of unfairness throughout the design and development of technical systems (Madaio et al., 2020). As a more technical operationalization, much fair ML research focuses on developing metrics and toolkits that practitioners can use to automatically measure and improve fairness in decision-making systems (Ojewale et al., 2025; Wong et al., 2023).
Fairness Challenges in Practice
Prior work has repeatedly shown that efforts to operationalize fairness within organizations are deeply challenging. Practitioners struggle to reconcile incompatible fairness metrics (Madaio et al., 2022; Moharana et al., 2025; Smith et al., 2025; Koenecke et al., 2023), work with messy real-world data (Cramer et al., 2019; Holstein et al., 2019; Rieke et al., 2022; Guerdan et al., 2025; Passi and Barocas, 2019; Sambasivan et al., 2021; Pang et al., 2023), and adapt general-purpose academic tools and metrics to their specific real-world needs (Deng et al., 2022; Harvey et al., 2025, 2024; Lee and Singh, 2021; Balayn et al., 2023; Kaur et al., 2020; Richardson et al., 2021; Ojewale et al., 2025; Madaio et al., 2024; Morley et al., 2020; Corpus et al., 2025). Challenges are exacerbated by organizational barriers to communicating and collaborating across different teams and within cross-functional teams (Deng et al., 2023b), or to incorporating the perspectives of multiple stakeholders, including end users, into fairness assessments and improvements (Deng et al., 2023a; Costanza-Chock et al., 2022; Madaio et al., 2020).
@@
With the notable exception of Smith et al. (Smith et al., 2023b, a, 2024, 2025), most prior work on implementing fairness in practice has focused broadly on general-purpose ML or specifically on fairness in classification models, NLP models, or other non-RS applications of ML. However, we expect the issues raised by prior work to be even more pressing in the RS space, where fewer academic toolkits are available to scaffold fairness efforts, and where the interests of multiple stakeholders (providers and users) may be in direct competition. In fact, Smith et al. have previously identified tensions between competing stakeholders through interview studies with individuals who interact with RS but do not work directly on technical RS teams, including providers (content creators and dating app users) (Smith et al., 2024) and ML fairness experts (Smith et al., 2025). Most similar to our work, Smith et al. have conducted interview studies with RS practitioners themselves: employees of an anonymous “commercial consumer-provider recommender system” (Smith et al., 2023a) and of a microlending platform (Smith et al., 2023b). Both studies explored challenges related to selecting fairness metrics in the context of RS, again identifying competing stakeholder interests as a key challenge (Smith et al., 2023b, a). In contrast, we interview technical RS practitioners across multiple big tech companies, and look beyond fairness metrics to how fairness is incorporated into RS workflows more broadly. In summary, we build on the extensive body of prior work by focusing explicitly on the RS workflow, with a spotlight on how RS practitioners across multiple big tech companies navigate not only technical challenges stemming from RS-specific complexity, but also organizational challenges stemming from cross-team interactions with legal, data, and fairness practitioners.